OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply. (9,3) (-3,-6) A. y - 6 = (x - 3) B. y - 3 = (x - 9) C. y + 6 = (x + 3) D. y - 3 = (x - 9) E. y + 6 = (x + 3) F. y - 6 = (x - 3)

OpenStudy (amistre64):

youve got point slope equations

OpenStudy (amistre64):

y - Py = (x-Px) is the format to watch for

OpenStudy (amistre64):

where P(x,y) indicates Px and Py

OpenStudy (anonymous):

so confuses

OpenStudy (amistre64):

on which part?

OpenStudy (anonymous):

the slope is 3/4 right?

OpenStudy (anonymous):

everything i dont know how to do it

OpenStudy (anonymous):

it is a online class i dont need to know how to do it

OpenStudy (amistre64):

apparently you do need to know how to do it, otherwise you wouldnt be asking the question :)

OpenStudy (amistre64):

( 9 , 3 ) -(-3,-6) ---------- 12,9; m = 9/12 = 3/4 yep. so either you got a typo in there or none of these are right

OpenStudy (anonymous):

yeah i thought so!! none of these are right

OpenStudy (anonymous):

it doesn't make since

OpenStudy (anonymous):

right - non e of your list fit the two points maybe u'd better check the original quaestion

OpenStudy (anonymous):

i copy and pasted. their are no typos

OpenStudy (anonymous):

Which of the following equations describes the line shown below? that is the question

OpenStudy (anonymous):

and the two points on the line are (9,3) and (-3,-6)?

OpenStudy (anonymous):

correct. yes

OpenStudy (anonymous):

i tried looking at your URL but it denied me

OpenStudy (anonymous):

here

OpenStudy (amistre64):

whoever typed up the question for the site missed it; try picking the ones that correspond to it as tho it had the right slope

OpenStudy (amistre64):

bcde might all apply; butyou do have duplicates in there

OpenStudy (anonymous):

the line through the two point has eqaution y - 9 = 3/4 (x-3) or 4y - 36 = 3x -9 that is 4y = 3x + 27

OpenStudy (anonymous):

you have 3 sets of duplicates

OpenStudy (anonymous):

i think i may just do B C D E

OpenStudy (anonymous):

or is that bad if i do thatt